OpenStudy (anonymous):

A classroom has a length of 20 ft. and a width of 30 ft. The floor will be covered in tiles that are 2 ft long and 3 ft wide. How many tiles are needed to fill the classroom?

OpenStudy (anonymous):

10.

OpenStudy (anonymous):

srry but that is wrong: the area of the class room is 20*30=600 the area of one tile is 2*3=6 the number of tiles need to cover the classroom is just the area of the classroom divided by the area of each individual tile: 600/6= 100 tiles